Solve the following pair of linear equations by the substitution method
0.2x + 0.3y = 1.3
0.4x + 0.5y = 2.3

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given,
0.2x + 0.3y = 1.3
and
0.4x + 0.5y = 2.3
are the two equations.
From 1st equation,
we get,
x = (1.3- 0.3y)/0.2 _________________(1)
Now,
substitute the value of x in the given second equation to get,
0.4(1.3-0.3y)/0.2 + 0.5y = 2.3
⇒2(1.3 – 0.3y) + 0.5y = 2.3
⇒ 2.6 - 0.6y + 0.5y = 2.3
⇒ 2.6 – 0.1 y = 2.3
⇒ 0.1 y = 0.3
⇒ y = 3
Now, substitute the value of y in equation (1),
we get,
x = (1.3-0.3(3))/0.2 = (1.3-0.9)/0.2 = 0.4/0.2 = 2
Therefore, x = 2 and y = 3 is required solution.

Thanks